UT 2010-Jul-14
==============
Program CL3, PI Kraus

Observers: Fabien, Rafael, PJ

CHARA/CLIMB
TEL       E1-E2-W1-S1
POP       1-4-4-1

TRIPLET   E1-E2-W1  (start-up)
BEAM      B1-B2-B3
REF CART  E2

Starting
LT 13 July 2010, 20:30 
Weather: clear, very windy, r0 = 3 to 5 cm

NIRO ALIGN on 7 Sco


cal HD143112
21:04 scanning for fringes
      comm problems
      -4400/24000
      clock problems

22:22 -4679/23000
      recording fringes on cal HD156365
      Vis:0.57/0.54/0.46
      Counts:80/60/70
      too lo

22:56 restarting with other cal and slower speed
      HD159743
      -4485/22884
      recording data at 250Hz/DEST
23:01 -4392/22982
      Vis:0.54/0.45/0.39
      Counts:150/100/150      
23:07 -4300/23074      
23:16 -4200/23400

      mwc275
      recording fringes at 250Hz/DEST
23:31 -4234/23490
      Vis:0.27/0.08/0.08      
      Counts:150/120/110
23:33 -4200/23590      
23:41 -4074/23780
      switching to cal HD164104
0:00 -3771/
0:03 /23597
0:05 -3792/23749
     recording fringes on HD164104
     Vis:0.54/0.50/0.56
     Counts:110/70/110

0:21 switching to MWC340
     cal194708
     W1 completely lost the delay offset and gives crazy offset (~ -15000000 um)
     restart of W1 server+gui
     Crash of W1 server
     Crash of CLIMB_1
     Crash of CLIMB server - restarted 
      
      W1 computer needed to be rebooted at the telescope

2:38 back running
     -1042/

3:13  -873 / 20242
      recording fringes on cal 194708 at 250Hz/DEST
      Vis:0.32/0.51/0.26
      Counts: 82/63/45
3:20 -828/20563
3:35 -700/20916
     recording fringes on MWC340 at 250Hz/DEST
3:50 -572/20920 ?
     -500/21637
4:03 -440/
     had to move carts, there will be light to remove during data analysis
4:12 -340/21900

4:28 cal HD199598
      -500/21244
      recording fringes at 250Hz/DEST
4:35 -438/21423            
      Vis: 0.32/0.48/0.25
      Counts: 180/135/100
4:54 -250/21794

Summary
-------
seeing was ok.
lots of small technical problems slowed us down. 
1 bracket on MWC275.
1 bracket on MWC340.